# Learning to Steer Markovian Agents under Model Uncertainty

### v1

@@ E.2.1 Special Case: PMD with Exact Advantage-Value
where we use $C_{h}(s_{h})$ and $C_{h}^{\prime}(s_{h})$ to denote some state-dependent but action-independent value. According to Lem. E.6, under the above steering reward design, the dynamics of ${\bm{\pi}}_{1},...,{\bm{\pi}}_{t},...,{\bm{\pi}}_{T}$ can be described by the following dynamics:
@@
The proof is a directly application of Lem. E.7. ∎
@@
Proof for Lem. E.9 and Thm. E.10
The proof is by directly applying Lem. E.7 and Thm. E.8 since NPG is a special case of PMD with KL-Divergence as Bregman Divergence. For any ${\bm{\pi}},{\widetilde{\bm{\pi}}}\in\Pi^{+}$, we consider the dual variables $\theta,\widetilde{\theta}$ such that:
@@
Choice of $U_{\max}$ in Lem. E.9
By applying Lem. E.7 and Thm. E.8, we consider the following choice of $U_{\max}$
@@
When ${\bm{\pi}}^{*}\in\Pi^{+}$, we can directly apply Thm. E.8 with ${\widetilde{\bm{\pi}}}\leftarrow{\bm{\pi}}^{*}$, and choosing $U_{\max}$ correspondingly following Eq. (8).
@@
E.2.2 The General Incentive Driven Agents under Assump. B
Theorem E.11 (Formal Version of Thm. 4.2 for the general PMD).
@@
Remark E.12.
In Thm. E.2.2, our bound for $U_{\max}$ here is just a worst-case bound to handle the noisy updates in the worst case. With high probability, the dual variable $\theta_{t}$ will converge to $\widetilde{\theta}$ and the steering reward does not have to be as large as $U_{\max}$.
@@
Here we choose $\nu^{n}_{t,h}(s_{h},a_{h}^{n},{\bm{\pi}}_{t}):=\frac{1}{\gamma}\cdot(\widetilde{\theta}^{n}_{h}(a_{h}^{n}|s_{h})-\theta^{n}_{t+1,h}(a_{h}^{n}|s_{h}))$, where $\widetilde{\theta}$ denotes the dual variable of policy $\widetilde{\bm{\pi}}$ and $\gamma$ will be determined later. Comparing with the design in the proof of Thm. E.8, here the “driven term” $\nu^{n}_{h}$ need to depend on ${\bm{\pi}}_{t}$ because of the randomness in updates.
@@
As we discuss in Example. E.3, Assump. C is satisfied with $\mu=1$. The proof is a direct application of Thm. E.8 with the same choice of dual variables as Eq. (7).
Appendix F Missing Proofs for Existence when the True Model $f^{*}$ is Unknown
@@
where we define $L_{M}(\tau_{T_{0}}):=\sum_{t=1}^{T_{0}}l_{M}({\bm{\pi}}_{t},{\bm{u}}_{t},{\bm{\pi}}_{t+1})$ and $L_{M}(\bar{\tau}_{T_{0}}):=\sum_{t=1}^{T_{0}}l_{M}({\bm{\pi}}_{t},{\bm{u}}_{t},\bar{{\bm{\pi}}}_{t+1})$. Besides, by applying Lem. F.1 on $l_{M}$ defined above and applying Markov inequality and the union bound over all $f\in\mathcal{F}$, we have $\Pr(\mathcal{E})\geq 1-\delta$. On the event $\mathcal{E}$, we have:
@@
Consider the steering strategy $\psi({\bm{\pi}})=u_{\bm{\pi}}$. Given any $f\in\mathcal{F}$, and the trajectory sampled by $\psi$ and $f$, by Lem. F.2, w.p. $1-\frac{\delta}{|\mathcal{F}|}$, we have:
@@
Appendix G Generalization to Partial Observation MDP Setup
G.1 POMDP Basics
@@
G.2 Steering Process as a POMDP
@@
Definition G.1 ($m$-Decodable POMDP).
@@
In our steering setting, if for any $f\in\mathcal{F}$, $f$ is $m$-decodable, we just need to learn a steering strategy $\psi:=(\mathcal{O}\times\mathcal{U})^{m}\times\mathcal{O}\rightarrow\mathcal{U}$, which predicts the steering reward given the past $m$-step history. This is the motivation for our experiment setup in the Grid World Stag Hunt game in Sec. 6.1. More concretely, we assume the agents trajectories in the past few steps can be used as sufficient statistics for the current policy, and use them as input of the steering strategy (see Appx. H.2.2 for more details).
@@
Appendix H Missing Experiment Details
H.1 About Initialization in Evaluation
@@
H.2 Experiments for Known Model Setting
H.2.1 Experiment Details in Normal-Form Stag Hunt Game
@@
Figure 5: Trade-off between Steering Gap (Left) and Steering Cost (Right). (averaged over 5x5 uniformly distributed grids as initializations of ${\bm{\pi}}_{1}$, see Appx. H.1).
H.2.2 Experiment Details in Grid-World Version of Stag Hunt Game
@@
H.2.3 Experiments in Matching Pennies
@@
We follow the same steering strategy and training setups for Stag Hunt Game in Appx. H.2.1. The agents follow NPG to update the policies with learning rate $\alpha=10$.
@@
Figure 6: Experiments in MatchingPennies. (a) $x$ and $y$ axes correspond to the probability that agents take Head. Black dots mark the initial policies, and red curves represents the trajectories of agents policies. The steering strategy to plot the figure is trained with $\beta=25$. (b) We compare $\beta=10,25,100$. Error bar shows 95% confidence intervals. (averaged over 5x5 uniformly distributed grids as initializations of ${\bm{\pi}}_{1}$, see Appx. H.1)
H.3 Experiments for Unknown Model Setting
H.3.1 Details for Experiments with Small Model Set $\mathcal{F}$
@@
The training of $\psi_{0.7}^{*}$ and $\psi_{1.0}^{*}$ follow the similar experiment setup as Appx. H.2.2, except here the agents adopt random learning rates. For the choice of $\beta$, we train the optimal steering strategy with $\beta\in\{10,20,30,40,50,60,70,80,90,100\}$ for both $f_{0.7}$ and $f_{1.0}$, and choose the minimal $\beta$ such that the resulting steering strategy can achieve almost 100% accuracy (i.e. $\Delta_{\psi,f}\leq\varepsilon$ for almost all 5x5 uniformly distributed initial policies generated by process in Appx. H.1). As we reported in the main text, we obtain $\beta=70$ for $f_{0.7}$ and $\beta=20$ for $f_{1.0}$.
@@
H.3.2 Details for Experiments with Large Model Set $\mathcal{F}$
We first highlight here that exploring and identifying $\lambda^{n}$ is necessary. Because the mediator can use the same incentive to drive those “opportunistic” agents with small $\lambda^{n}$ for a larger update step towards the desired policy, comparing with those “steady” ones with larger $\lambda^{n}$.
We set $U_{\max}=2.0$, and the random exploration strategy (red curve in the left sub-plot in Fig. 3) will sample the steering reward uniformly from the interval $[0,U_{\max}]$. We use the PPO (Raffin et al.,, 2021) to train of exploration policy and also the steering strategy given hidden model.
@@ H.3.2 Details for Experiments with Large Model Set $\mathcal{F}$
$\psi^{\text{Explore}}\leftarrow\arg\max_{\psi}\frac{1}{|\mathcal{F}|}\sum_{f\in\mathcal{F}}\mathbb{E}_{\psi,f}[\sum_{n\in[N]}\Pr(\lambda^{n}|{\bm{\pi}}_{1},u_{1},...,{\bm{\pi}}_{\widetilde{T}},u_{\widetilde{T}}{\bm{\pi}}_{\widetilde{T}+1})].$
@@
To plot the results in the middle and right sub-plots in Fig. 3, for each model $f^{*}\in\{f_{A},f_{B},f_{C},f_{D}\}$, we train two steering strategies (with the same state design in Eq. (11)). The first one is the oracle strategy, which starts with ${\bm{\pi}}_{1}$ and steering for $T=500$ steps. The second one is the exploitation policy paired with the exploration policy. Because we do not know what ${\bm{\pi}}_{\widetilde{T}}$ is after executing the exploration policy for $\widetilde{T}$ steps. During the training of the exploitation policy, we use random initialization, and set $T=470$, which subtract the exploration steps.
During the evaluation, for FETE, we steer with the exploration policy for the first 30 steps, and execute the exploitation policy for the rest. The results are averaged over 5 seeds and two-sigma error bar is shown.
H.4 A Summary of the Compute Resources by Experiments in this Paper

### v2

@@ D.2.1 Special Case: PMD with Exact Advantage-Value
where we use $C_{h}(s_{h})$ and $C_{h}^{\prime}(s_{h})$ to denote some state-dependent but action-independent value. According to Lem. D.6, under the above steering reward design, the dynamics of ${\bm{\pi}}_{1},...,{\bm{\pi}}_{t},...,{\bm{\pi}}_{T}$ can be described by the following dynamics:
@@
The proof is a directly application of Lem. D.7. ∎
@@
Proof for Lem. D.9 and Thm. D.10
The proof is by directly applying Lem. D.7 and Thm. D.8 since NPG is a special case of PMD with KL-Divergence as Bregman Divergence. For any ${\bm{\pi}},{\widetilde{\bm{\pi}}}\in\Pi^{+}$, we consider the dual variables $\theta,\widetilde{\theta}$ such that:
@@
Choice of $U_{\max}$ in Lem. D.9
By applying Lem. D.7 and Thm. D.8, we consider the following choice of $U_{\max}$
@@
When ${\bm{\pi}}^{*}\in\Pi^{+}$, we can directly apply Thm. D.8 with ${\widetilde{\bm{\pi}}}\leftarrow{\bm{\pi}}^{*}$, and choosing $U_{\max}$ correspondingly following Eq. (8).
@@
D.2.2 The General Incentive Driven Agents under Assump. B
Theorem D.11 (Formal Version of Thm. 4.2 for the general PMD).
@@
Remark D.12.
In Thm. D.2.2, our bound for $U_{\max}$ here is just a worst-case bound to handle the noisy updates in the worst case. With high probability, the dual variable $\theta_{t}$ will converge to $\widetilde{\theta}$ and the steering reward does not have to be as large as $U_{\max}$.
@@
Here we choose $\nu^{n}_{t,h}(s_{h},a_{h}^{n},{\bm{\pi}}_{t}):=\frac{1}{\gamma}\cdot(\widetilde{\theta}^{n}_{h}(a_{h}^{n}|s_{h})-\theta^{n}_{t+1,h}(a_{h}^{n}|s_{h}))$, where $\widetilde{\theta}$ denotes the dual variable of policy $\widetilde{\bm{\pi}}$ and $\gamma$ will be determined later. Comparing with the design in the proof of Thm. D.8, here the “driven term” $\nu^{n}_{h}$ need to depend on ${\bm{\pi}}_{t}$ because of the randomness in updates.
@@
As we discuss in Example. D.3, Assump. C is satisfied with $\mu=1$. The proof is a direct application of Thm. D.8 with the same choice of dual variables as Eq. (7).
Appendix E Missing Proofs for Existence when the True Model $f^{*}$ is Unknown
@@
where we define $L_{M}(\tau_{T_{0}}):=\sum_{t=1}^{T_{0}}l_{M}({\bm{\pi}}_{t},{\bm{u}}_{t},{\bm{\pi}}_{t+1})$ and $L_{M}(\bar{\tau}_{T_{0}}):=\sum_{t=1}^{T_{0}}l_{M}({\bm{\pi}}_{t},{\bm{u}}_{t},\bar{{\bm{\pi}}}_{t+1})$. Besides, by applying Lem. E.1 on $l_{M}$ defined above and applying Markov inequality and the union bound over all $f\in\mathcal{F}$, we have $\Pr(\mathcal{E})\geq 1-\delta$. On the event $\mathcal{E}$, we have:
@@
Consider the steering strategy $\psi({\bm{\pi}})=u_{\bm{\pi}}$. Given any $f\in\mathcal{F}$, and the trajectory sampled by $\psi$ and $f$, by Lem. E.2, w.p. $1-\frac{\delta}{|\mathcal{F}|}$, we have:
@@
Appendix F Generalization to Partial Observation MDP Setup
F.1 POMDP Basics
@@
F.2 Steering Process as a POMDP
@@
Definition F.1 ($m$-Decodable POMDP).
@@
In our steering setting, if for any $f\in\mathcal{F}$, $f$ is $m$-decodable, we just need to learn a steering strategy $\psi:=(\mathcal{O}\times\mathcal{U})^{m}\times\mathcal{O}\rightarrow\mathcal{U}$, which predicts the steering reward given the past $m$-step history. This is the motivation for our experiment setup in the Grid World Stag Hunt game in Sec. 6.1. More concretely, we assume the agents trajectories in the past few steps can be used as sufficient statistics for the current policy, and use them as input of the steering strategy (see Appx. G.2.2 for more details).
@@
Appendix G Missing Experiment Details
G.1 About Initialization in Evaluation
@@
G.2 Experiments for Known Model Setting
G.2.1 Experiment Details in Normal-Form Stag Hunt Game
@@
Figure 5: Trade-off between Steering Gap (Left) and Steering Cost (Right). (averaged over 5x5 uniformly distributed grids as initializations of ${\bm{\pi}}_{1}$, see Appx. G.1).
G.2.2 Experiment Details in Grid-World Version of Stag Hunt Game
@@
G.2.3 Experiments in Matching Pennies
@@
We follow the same steering strategy and training setups for Stag Hunt Game in Appx. G.2.1. The agents follow NPG to update the policies with learning rate $\alpha=10$.
@@
Figure 6: Experiments in MatchingPennies. (a) $x$ and $y$ axes correspond to the probability that agents take Head. Black dots mark the initial policies, and red curves represents the trajectories of agents policies. The steering strategy to plot the figure is trained with $\beta=25$. (b) We compare $\beta=10,25,100$. Error bar shows 95% confidence intervals. (averaged over 5x5 uniformly distributed grids as initializations of ${\bm{\pi}}_{1}$, see Appx. G.1)
G.3 Experiments for Unknown Model Setting
G.3.1 Details for Experiments with Small Model Set $\mathcal{F}$
@@
The training of $\psi_{0.7}^{*}$ and $\psi_{1.0}^{*}$ follow the similar experiment setup as Appx. G.2.2, except here the agents adopt random learning rates. For the choice of $\beta$, we train the optimal steering strategy with $\beta\in\{10,20,30,40,50,60,70,80,90,100\}$ for both $f_{0.7}$ and $f_{1.0}$, and choose the minimal $\beta$ such that the resulting steering strategy can achieve almost 100% accuracy (i.e. $\Delta_{\psi,f}\leq\varepsilon$ for almost all 5x5 uniformly distributed initial policies generated by process in Appx. G.1). As we reported in the main text, we obtain $\beta=70$ for $f_{0.7}$ and $\beta=20$ for $f_{1.0}$.
@@
G.3.2 Details for Experiments with Large Model Set $\mathcal{F}$
We set $U_{\max}=1.0$, and the random exploration strategy (red curve in the left sub-plot in Fig. 3) will sample the steering reward uniformly from the interval $[0,U_{\max}]$. We use the PPO (Raffin et al.,, 2021) to train of exploration policy and also the steering strategy given hidden model. To amplify the exploration challenge, we set $\beta^{n}=1$ when $\pi^{n}(\texttt{A})\leq 0.5$ and increase to $\beta^{n}=10$ when $\pi^{n}(\texttt{A})\leq 0.5$. As a result, if the mediator follows first-explore-then-exploit strategy and fail to distinguish avaricious agents from the normal ones, adopting large steering reward can lead to much worse performance.
@@ G.3.2 Details for Experiments with Large Model Set $\mathcal{F}$
$\psi^{\text{Explore}}\leftarrow\arg\max_{\psi}\frac{1}{|\mathcal{F}|}\sum_{f\in\mathcal{F}}\mathbb{E}_{\psi,f}[\sum_{n\in[N]}\Pr(\lambda^{n}|{\bm{\pi}}_{1},u_{1},...,{\bm{\pi}}_{\widetilde{T}},u_{\widetilde{T}}{\bm{\pi}}_{\widetilde{T}+1})].$ (13)
@@
To plot the results in the middle and right sub-plots in Fig. 3, for each model $f^{*}\in\{f_{1},f_{2},f_{3}\}$, we train three steering strategies (with the same state design in Eq. (11)). The first one is the oracle strategy, which starts with ${\bm{\pi}}_{1}$ and steering for $T=500$ steps. The second one is the FETE strategy, including an exploration policy and another exploitation policy. The exploration policy is trained following the above Eq. (13) with exploration horizon $\widetilde{T}=30$. Then, we estimate the model from samples generated by the exploration policy, and train another exploitation policy following the remaining step of FETE with exploitation horizon $T=470$ (Procedure 2). The third strategy FETE-RE is the same as FETE except we just use random policy as the exploration policy, and estimate the model by interaction samples generated by that. During the evaluation, for FETE and FETE-RE, we steer with the exploration policy for the first 30 steps, and execute the exploitation policy for the rest. The results are averaged over 5 seeds and two-sigma error bar is shown.
G.4 Explanation of the Consistency of the Adaption
We first want to highlight it is not easy to have an apples-to-apples comparison with (Canyakmaz et al.,, 2024). First, because the experiment setting in (Canyakmaz et al.,, 2024) does not present significant exploration challenges, we design and decide to evaluate both methods in the “avaricious agents” setting (Fig. 3). Second, SIAR-MPC is specialized for polynomial function classes and the dynamics tractable by MPC, making it difficult to generalize beyond that setting. Therefore, we have to do some necessary adaption. To ensure the fair comparison, we consider to use FETE-RE as the adaption of SIAR-MPC in our setting. For the exploration stage, FETE-RE aligns with SIAR-MPC in using random exploration. For the model identification and exploitation phase, FETE-RE adopts MLE estimation and RL methods to train the exploitation policy, which inherit the same inspirits as SIAR-MPC and also aligned with our original FETE.
From another perspective, the main focus of our empirical comparison between FETE and (Canyakmaz et al.,, 2024) is the impact of different exploration strategies on the final steering gap and steering costs. This is reasonable. Because from the discussion in Sec. 5.2, we can conclude that our FETE is more general compared with SIAR-MPC in (Canyakmaz et al.,, 2024) in terms of both the model estimation strategy and exploitation strategy. The main distinction and improvement of our FETE compared with (Canyakmaz et al.,, 2024) is our strategic exploration strategy.
G.5 A Summary of the Compute Resources by Experiments in this Paper
@@
Appendix H Discussion on Limitations and Societal Impact
Discussion on Limitations
Discussion on Societal Impact